OpenStudy (anonymous):

Express the complex number in trigonometric form. -6 + 6√3 i know to get r its √a^2+b^2 which is √36+324, so √360... but now what????????

OpenStudy (irishboy123):

\(-6 + \color{red}i \;6\sqrt{3} = 12(-\dfrac{1}{2} + \dfrac{\sqrt{3}}{2})\)

OpenStudy (anonymous):

12 (cos2pi/3 + sin2pi/3) ????